OpenStudy (anonymous):

You place a cup of 205F coffee on a table in a room that is 72F, and 10 minutes later, it is 195F. Approximately how long will it be before the coffee is 180F?

OpenStudy (anonymous):

there are a couple ways to do this but either way works with the difference in the temperatures, not the temperatures in other words work with \(205-72=133\) and \(195-72=123\)

OpenStudy (anonymous):

the easy way is to say that since it takes 10 minutes for the temp difference to go from 133 to 123 then you can model it as \[T=133\times \left(\frac{123}{133}\right)^{\frac{t}{10}}\]

OpenStudy (anonymous):

when the coffee is \(180\) degrees then the temperature difference is \(180-72=108\) and you can solve by setting \[108=133\times \left(\frac{123}{133}\right)^{\frac{t}{10}}\] and solve for \(t\)

OpenStudy (anonymous):

im so bad at math.. how would i get t?

OpenStudy (anonymous):

the other way is to use \[T=T_0e^{kt}\] which may be what you are taught, but then you have to solve for \(k\) first is that what you are supposed to do? or are you just supposed to get the answer?

OpenStudy (anonymous):

25 mins!

OpenStudy (anonymous):

i got it thank you.

OpenStudy (anonymous):

ok good, but i get 26.6 about, if you are just approximating then 25 is close

OpenStudy (anonymous):

yw